OpenStudy (anonymous):

What is the slope and y-intercept of the equation 3(y − 2) + 6(x + 1) − 2 = 0

jigglypuff314 (jigglypuff314):

the slope is -2 the y-intercept is 2/3 because 3(y-2)+6(x+1)-2=0 can be factored out to 3y-6+6x+6-2=0 which can then be reduced to 3y+6x-2=0 then put into slope-intercept form y=-2x+(2/3)